# Continual Lie algebras determined by chain complexes

### v5

@@ 3.3. Double cochain complex: Godbillon-Vey type example
In this subsection we provide the explicit example for Theorem (1) in the case of the orthogonality condition (2.5) applied to the particular case when, in the consideration of the previous subsection, $\Phi=\chi$. In particular, in differential geometry, the case of a foliation $\mathcal{F}$ of codimension one defined by a one-form on a three-dimensional manifold, and the formulation of the Godbillon-Vey cohomology class, are included in this consideration.
@@
Thus, for $\alpha_{1}^{(R)}\in C^{n^{\prime},m^{\prime}}(\mathcal{F})$ one has
@@
and $n+1=n+n^{\prime}-r$, $m+1=m+m^{\prime}-t$, and (3.9) is possible only when $n^{\prime}=r+1$, $m^{\prime}=t+1$, $0\leq r\leq n$, $0\leq t\leq n$, and $\alpha_{1}^{(R)}\in C^{r+1,t+1}(\mathcal{F})$. If we require from (3.8) that for $\alpha_{1}^{(L)}\in C^{\alpha,\beta}(\mathcal{F})$, $\chi=\alpha_{1}^{(L)}\cdot\delta^{n,m}\chi$, then $n=\alpha+n+1-r^{\prime}$, $\alpha=r^{\prime}-1$, i.e., $\alpha$ is smaller than the common degree which is not possible and thus such $\alpha_{1}^{(L)}$ does not exist.
Then, as a result of (2.2), we obtain the system of differential relations:
$0=\chi\cdot\delta^{n,m}\chi,\qquad(1),\quad 0=\left(\delta^{n,m}\chi\right)\cdot\delta^{n,m}\chi,\qquad(3)$
$\delta^{n,m}\chi=\chi\cdot\alpha^{(R)}_{1},\quad 0=\delta^{n,m}\chi\cdot\alpha^{(R)}_{1}+(-1)^{n}\chi\cdot\delta^{r+1,t+1}\alpha^{(R)}_{1},$ (3.10)
@@
and the rest of the system (2.2) collapses since further its branches follow from $(3)$ which is trivial.
@@
Let us denote ${\bf h}_{n}=\left(h_{1},\ldots,h_{n}\right)$, an $n$-tuple of holonomy mappings (see subsection 3.1). In this case, for forms $\left\{\chi,\;\delta^{n,m}\chi,\;\alpha^{(R)}_{1},\;\delta^{r+1,t+1}\alpha_{1}^{(L)}\right\}$, define the following continual Lie algebra by identifying the differential forms with generators of $\mathcal{G}(\mathcal{F})$ as
$X_{+}=\chi,\quad X_{-}=\delta^{n,m}\chi,\quad H=\alpha^{(R)}_{1},\quad H^{*}=\delta^{r+1,t+1}\alpha^{(R)}_{1},$ (3.11)
and the commutation relations (in a non-principal grading):
$\left[X_{+}({\bf h}_{n}),X_{-}({\bf h^{\prime}}_{n})\right]$ $=$ $0,\quad\left[X_{+}({\bf h}_{n}),H({\bf h}_{r+1})\right]=X_{-}\left(K_{+1,0}\left({\bf h}_{n},{\bf h}_{r+1}\right)\right),$
$\left[X_{-}({\bf h}_{n+1}),H({\bf h}_{r+1})\right]$ $+$ $(-1)^{n}\left[X_{+}({\bf h}_{n}),H^{*}({\bf h}_{r+2})\right]=0.$ (3.12)
@@
Note that an element of $C^{n,m}(\mathcal{F})$ is an $n$-form $\omega(h_{1},\ldots,h_{n})$ depending on $n$ holonomy maps. Thus the space of continual roots is provided by the space of holonomy embeddings (see subsection 3.1). Taking into account (3.3), we find the kernels
$K_{+1,-1}\left({\bf h}_{n},{\bf h}_{r+1}\right)$ $=$ $0,\quad K_{+1,0}\left({\bf h}_{n},{\bf h}_{r+1}\right)={\bf h}_{n+1},$
$K_{0,-1}\left({\bf h}_{n},{\bf h}_{r+1}\right)$ $=$ ${\bf h}_{n+2},\quad K_{0,1}\left({\bf h}_{n},{\bf h}_{r+1}\right)={\bf h}_{n+2}.$
@@
It is easy to check that the Jacobi identity for generators (3.11) are fulfilled. In the case of a codimension one foliation defined by a one-form, $\chi$, $\alpha^{(R)}_{1}$, $\chi\in C^{1,m}(\mathcal{F})$. Recall [10] that the Godbillon-Vey cohomology class is given by $\left[\alpha^{(R)}_{1}\wedge\delta^{1,m}\alpha^{(R)}_{1}\right]$. The construction above clarifies the Lie-algebraic meaning of this cohomology class.
@@ 4. Conclusion
In conclusion, we would like to mention a few directions of development and further applications of the material presented in this paper. We propose a way to associate a continual Lie algebra to a chain complex. Thus the properties, in particular, the Jacobi identity, kernels, and relations of resulting continual Lie algebras depend on the kind of chain complex spaces as well as on the set of their parameters. One can think of introducing various types of products suitable for the construction of systems of differential relations more complicated than (2.2). In our particular case (Section 2), in order to make connection with continual Lie algebras, we have chosen the commutator (2.17) (with respect to the original product defined on bicomplex spaces) as the simplest natural product. One could think of other possibilities which would be coherent with the orthogonality condition (2.5).
In our exposition, the standard form of chain complexes was involved. Nevertheless, one can consider more complicated setups, in particular, chain complexes where differentials act in non-trivial ways with respect to indices of spaces (c.f. examples in [11]). That would lead to alternative forms of systems of differential relations as well as of compatibility conditions. What could be especially interesting, is to treat multiple chain complexes containing combinations of a few chain-cochains.
The example of the Čhech-de Rham cohomology that we study in Section 3 comes from the differential geometry of foliations. In classics, the orthogonality condition applied to elements and their differentials of one particular bicomplex space, boils down to the integrability condition, and leads to the Frobenius theorem. Then it delivers the Godbillon-Vey cohomology class whose geometric meaning is not yet completely studied [9]. As for further applications in differential geometry, starting from the orthogonality condition, it would be interesting to find other cohomology invariants related to the Čech-de Rham bicomplex for foliations, and to understand their geometric meaning. The constructions considered in this paper can be also used for the cohomology theory of smooth manifolds, in particular, in various approaches to the construction of cohomology classes (cf., in particular, [13]). Wide applications are awaiting for new examples of continual Lie algebras in the field of integrable models [15, 14]. The cases of non-commutative fields used to define continual Lie algebras would also be useful for in non-commutative geometry.
@@ Acknowledgment
The author is supported by the Institute of Mathematics, Academy of Sciences of the Czech Republic (RVO 67985840).

### v6

@@ 3.3. Double cochain complex: Godbillon-Vey type example
In this subsection we provide the explicit example for Theorem (LABEL:mainpro) in the case of the orthogonality condition (LABEL:ortho) applied to the particular case when, in the consideration of the previous subsection, $\Phi=\chi$. In particular, in differential geometry, the case of a foliation $\mathcal{F}$ of codimension one defined by a one-form on a three-dimensional manifold, and the formulation of the Godbillon-Vey cohomology class, are included in this consideration.
@@
First, let us check the condition for (3.8). Since ${\bf h}_{n}$ and ${\bf h}_{n+1}$ has to have the same ${\bf h}_{n}$, (according to the derivation of chain condition), thus n+(n+1)r=0n+(n+1)-r=0, but $r=n$, thus, n+(n+1)n=0n+(n+1)-n=0, i.e., n=1n=-1 which is not possible. Even if ${\bf h}_{n}$ and ${\bf h}_{n+1}$ has not to have the same part, we obtain n+(n+1)r=0n+(n+1)-r=0 $r=2n+1$, i.e., $r>n$ which is not also possible. Thus, for $\alpha_{1}^{(R)}\in C^{n^{\prime},m^{\prime}}(\mathcal{F})$ one has
@@
and $n+1=n+n^{\prime}-r$, $m+1=m+m^{\prime}-t$, and (3.9) is possible only when $n^{\prime}=r+1$, $m^{\prime}=t+1$, $0\leq r\leq n$, $0\leq t\leq n$, and $\alpha_{1}^{(R)}\in C^{r+1,t+1}(\mathcal{F})$. If we require from (3.8) that for $\alpha_{1}^{(L)}\in C^{\alpha,\beta}(\mathcal{F})$, $\chi=\alpha_{1}^{(L)}\cdot\delta^{n,m}\chi$, then $n=\alpha+n+1-r^{\prime}$, $\alpha=r^{\prime}-1$, i.e., $\alpha$ is smaller than the common degree which is not possible and thus such $\alpha_{1}^{(L)}$ does not exist. Then, as a result of (2.2), we obtain the system of relations:
$0=\chi\cdot\delta^{n,m}\chi,\qquad(1),\quad 0=\left(\delta^{n,m}\chi\right)\cdot\delta^{n,m}\chi\quad(3),\quad\delta^{n,m}\chi=\chi\cdot\alpha^{(R)}_{1}$, (3.10)
@@
and other relations of the system (2.2) at this level collapse since further its branches follow from $(3)$ which is trivial.
@@
Let us denote ${\bf h}_{n}=\left(h_{1},\ldots,h_{n}\right)$, an $n$-tuple of holonomy mappings (see subsection ??) (the space of parameters, here given by holonomy data). In this example, for forms $\left\{\chi({\bf h}_{n}),\;\delta^{n,m}\chi({\bf h}^{\prime}_{n}),\;\alpha^{(R)}_{1}({\bf h}^{\prime\prime}_{r+1})\right\}$ define the local part of the following continual Lie algebra by identifying the differential forms with generators of $\mathcal{G}(\mathcal{F})$ as
$X_{+1}({\bf h}_{n})=\chi({\bf h}_{n}),\;X_{-1}({\bf h}^{\prime}_{n+1})=\delta^{n,m}\chi({\bf h}^{\prime}_{n}),\;X_{0}({\bf h}^{\prime\prime}_{r+1})=\alpha^{(R)}_{1}({\bf h}^{\prime\prime}_{r+1}),$ (3.11)
The next step relation $0=\delta^{n,m}\chi\cdot\alpha^{(R)}_{1}+\gamma_{n,r+1}\;\chi\cdot\delta^{r+1,t+1}\alpha^{(R)}_{1}$, in the system (2.2) gives rise a higher grading generator identified with $\delta^{r+1,t+1}\alpha^{(R)}_{1}({\bf h}^{\prime\prime}_{r+1})$.
According to our derivation of Section ??, though $n$ and $m$ are fixed in the (3.11), the actual tuples ${\bf h}_{n}$, ${\bf h}^{\prime}_{n}$ in $\chi({\bf h}_{n})$ and $\delta^{n,m}\chi({\bf h}^{\prime}_{n})$ of (3.8) could be different with the same $n$.
Substituting in (3.10) the product $\cdot$ with the graded commutator, we obtain the following commutation relations in the principal grading:
$\left[X_{+1}({\bf h}_{n}),X_{0}({\bf h^{\prime}}_{n+1})\right]=0,\quad\left[X_{0}({\bf h}_{n+1}),X_{0}({\bf h}^{\prime}_{n+1})\right]=0,$
$\left[X_{+1}({\bf h}_{n}),X_{-1}(\widetilde{\bf h}_{r+1})\right]=X_{0}\left({\bf h}^{\prime}_{n+1}\right),$
@@
where we skip $gr$ in notation of commutators here and what follows. Note also, that the product of two identical differential forms of [5] vanishes. Therefore, $\chi({\bf h}_{n}),\chi({\bf h}^{\prime}_{n})]=0$, i.e., $X_{+1}({\bf h}_{n}),X_{+1}({\bf h}_{n})]=0$, and $X_{0}(\widetilde{\bf h}_{r+1}),X_{0}(\widetilde{\bf h}^{\prime}_{r+1})]=0$. In this example we use a non-standard grading. Note that an element of $C^{n,m}(\mathcal{F})$ is an $n$-form $\omega(h_{1},\ldots,h_{n})$ depending on $n$ holonomy maps. Thus, the space of continual roots is provided by the space of holonomy embeddings (see subsection ??). From (3.3) we extract directly the kernels
$K_{+1,0}\left({\bf h}_{n},{\bf h}^{\prime}_{r+1}\right)=0,\quad K_{0,0}\left({\bf h}_{n},{\bf h}^{\prime}_{r+1}\right)=0,\quad K_{+1,-1}\left({\bf h}_{n},\widetilde{\bf h}_{r+1}\right)={\bf h}^{\prime}_{n+1},$
@@
Therefore, $K_{+1,+1}({\bf h}_{n},{\bf h}^{\prime}_{n})=0$ and $K_{0,0}({\bf h}^{\prime\prime}_{r+1},{\bf h}^{\prime\prime}_{r+1})=0$. In the case of a codimension one foliation defined by a one-form, $\chi$, $\alpha^{(R)}_{1}$, $\chi\in C^{1,m}(\mathcal{F})$. Recall [10] that the Godbillon-Vey cohomology class is given by $\left[\alpha^{(R)}_{1}\wedge\delta^{1,m}\alpha^{(R)}_{1}\right$]. The construction above clarifies the Lie-algebraic meaning of this cohomology class.
We then check the graded the Jacobi identity for generators (3.11). For instance, let us do that for the generators $X_{+1}({\bf h}_{n})$, $X_{-1}(\widetilde{\bf h}_{r+1})$, and $X_{0}\left({\bf h}^{\prime}_{n+1}\right)$. From (3.3) we obtain
$\left[X_{-1}(\widetilde{\bf h}_{n+1}),X_{0}({\bf h}^{\prime}_{r+1})\right]=-\gamma_{n,r+1}\left[X_{+1}(\widetilde{\bf h}_{n}),Y_{0}({\bf h}^{\prime}_{r+2})\right].$
Usiing the formula for the double graded commutator, we see that
$\left[X_{+1}({\bf h}_{n}),\left[X_{+1}({\bf h}_{n}),Y_{0}({\bf h}^{\prime}_{r+2})\right]\right]$
$=\left[\left[X_{+1}({\bf h}_{n}),X_{+1}({\bf h}_{n})\right],Y_{0}({\bf h}^{\prime}_{r+2})\right]+(-1)^{nn}\left[X_{+1}({\bf h}_{n}),\left[X_{+1}({\bf h}_{n}),Y_{0}({\bf h}^{\prime}_{r+2})\right]\right],$
when $n^{2}=2k+1$, $k\in\mathbb{Z}$, $\left[X_{+1}({\bf h}_{n}),\left[X_{-1}(\widetilde{\bf h}_{r+1}),X_{0}\left({\bf h}^{\prime}_{n+1}\right)\right]\right]=0$. Then it follows
$\gamma_{n+1,r+1}\;\left[X_{0}({\bf h^{\prime}}_{n+1}),\left[X_{+1}({\bf h}_{n}),X_{-1}(\widetilde{\bf h}_{r+1})\right]\right]$
$+\gamma_{n,n+1}\;\left[X_{+1}({\bf h}_{n}),\left[X_{-1}(\widetilde{\bf h}_{r+1}),X_{0}({\bf h^{\prime}}_{n+1})\right]\right]$
$+\gamma_{r+1,n}\;\left[X_{-1}(\widetilde{\bf h}_{r+1}),\left[X_{0}({\bf h^{\prime}}_{n+1}),X_{+1}({\bf h}_{n})\right]\right]$
$=\gamma_{n+1,r+1}\left[X_{0}({\bf h^{\prime}}_{n+1}),X_{0}({\bf h^{\prime}}_{n+1})\right]+\gamma_{n,n+1}\gamma_{n,r+1}\left[X_{+1}({\bf h}_{n}),\left[X_{+1}({\bf h}_{n}),Y_{0}({\bf h}^{\prime}_{r+2})\right]\right]$
$+\gamma_{r+1,n}\;\left[X_{-1}(\widetilde{\bf h}_{r+1}),0\right]=0,$
and the Jacobi identity is satisfied. Thus, the generators defined in (3.11) satisfy the Jacobi identity, confirming the consistency of the continual Lie algebra $\mathcal{G}(\mathcal{F})$ associated with the chain complex.
@@ 4. Conclusion
In conclusion, we would like to mention a few directions of development and further applications of the material presented in this paper. We propose a way to associate a continual Lie algebra to a chain complex. Thus the properties, in particular, the Jacobi identity, kernels, and relations of resulting continual Lie algebras depend on the kind of chain complex spaces as well as on the set of their parameters. One can think of introducing various types of products suitable for the construction of systems of relations more complicated than (2.2). In our particular case (Section ??), in order to make connection with continual Lie algebras, we have chosen the commutator (2.17) (with respect to the original product defined on bicomplex spaces) as the simplest natural product. One could think of other possibilities which would be coherent with the orthogonality condition (LABEL:ortho).
In our exposition, the standard form of chain complexes was involved. Nevertheless, one can consider more complicated setups, in particular, chain complexes where differentials act in non-trivial ways with respect to indices of spaces (c.f. examples in [11]). That would lead to alternative forms of systems of relations as well as of compatibility conditions. What could be especially interesting, is to treat multiple chain complexes containing combinations of a few chain-cochains.
The example of the Čhech-de Rham cohomology that we study in Section ?? comes from the differential geometry of foliations. In classics, the orthogonality condition applied to elements and their differentials of one particular bicomplex space, boils down to the integrability condition, and leads to the Frobenius theorem. Then it delivers the Godbillon-Vey cohomology class whose geometric meaning is not yet completely studied [9]. As for further applications in differential geometry, starting from the orthogonality condition, it would be interesting to find other cohomology invariants related to the Čech-de Rham bicomplex for foliations, and to understand their geometric meaning. The constructions considered in this paper can be also used for the cohomology theory of smooth manifolds, in particular, in various approaches to the construction of cohomology classes (cf., in particular, [13]). Wide applications are awaiting for new examples of continual Lie algebras in the field of integrable models [15, 14]. The cases of non-commutative fields used to define continual Lie algebras would also be useful for in non-commutative geometry.
@@ Acknowledgment
The author is supported by the Institute of Mathematics, Academy of Sciences of the Czech Republic (RVO 67985840). The author would like to thank an unknown referee for valuable corrections and comments.